# Sound Source Localization for Spatial Mapping of Surgical Actions in Dynamic Scenes

### v1

@@ 1 Introduction
Surgical procedures are inherently dynamic, involving complex interactions among surgical staff, patients, instruments, and other medical devices. Understanding these interactions — referred to as surgical scene understanding — will be a cornerstone of next-generation computer-aided surgery and a crucial enabler for intelligent and autonomous surgical systems capable of assisting or operating independently.
In surgical scene understanding research, a body of previous work proposed to explicitly model the relationships among surgeons, patients, and instruments using semantic and spatial information, e.g., as action triplets [1] or scene graphs [2], and use these intermediate representations for downstream tasks such as surgical phase or role prediction. In contrast to explicit modeling approaches, end-to-end training has recently demonstrated strong performance in predicting high-level surgical concepts (e.g., surgical phases) from low-level observations (e.g., individual video frames) [3, 4]. However, end-to-end learning-based approaches can be prone to learning non-causal relationships and dataset-specific biases [5]. Constructing a fine-grain digital representation of the surgical scene, often called a surgical digital twin, can be a promising framework to address this limitation. This resulting representation can then be used for geometric scene understanding in a subsequent stage, effectively decoupling low-level data processing from high-level semantic inference [5, 6]. While current approaches for the generation of digital surgery representations predominantly rely on visual data from optical camera systems, the integration of multimodal information has been shown to be beneficial for surgical scene understanding in end-to-end learning frameworks [3, 4] and is expected to play a pivotal role in advancing next-generation data-driven computer-aided surgery systems [7]. In this context, the integration of multimodal data in digital surgical scene representations has great potential to achieve a more fine-grain modeling of surgical procedures.
Recently, acoustic events generated during tool–tissue interactions—such as sawing, chiseling, or drilling—have been shown to provide valuable insights into surgical tasks [8, 9, 10, 11]. Furthermore, in our own previous work, we proposed a proof-of-concept for the localization of acoustic events in surgery based on a microphone array and acoustic beamforming in 2D image space [12]. The integration of acoustic information into digital surgical scene representations, however, requires spatial and temporal awareness to enable effective surgical scene understanding.
@@
  • We propose a novel concept to generate 4D audio-visual representations of surgical scenes by fusing dynamic point clouds obtained using a RGB-D camera with acoustic localization information obtained from an phased microphone array, also known as acoustic camera.
@@
To the best of the author’s knowledge, no comparable approach currently exists within the state of the art. The code and associated data will be made publicly available [^1] The code is available under https://github.com/matthiasseibold/ssl_v2. The associated data will be made available upon acceptance..
@@ 2 Related Work
To establish a systematic framework for describing and analyzing surgical procedures, previous research has introduced methods at different levels of abstraction. Examples include hierarchical context models of surgical tasks and sub-tasks within Surgical Process Models [13], as well as automated extraction of surgical phases from intraoperative video recordings [14]. Automated analysis and description of surgical workflows can support workflow optimization, facilitate automated documentation for reporting and education, detect complications or deviations from preoperative plans, and enable real-time alerts in the event of adverse occurrences [15].
More recently, action triplets have been proposed as an extension of high-level surgical phase models, explicitly capturing the interactions inherent to surgical procedures [1]. Furthermore, graph-based representations have been introduced that model surgical scenes as structured graphs, where nodes correspond to relevant actors and objects in the operating room and edges represent spatial and functional relationships between them [2, 16, 17]. State-of-the-art approaches in this area primarily rely on visual data acquired from RGB or RGB-D cameras to derive both semantic and spatial information for modeling the surgical context. While end-to-end learning-based approaches achieve state-of-the-art performance for surgical scene graph prediction and leverage multimodal information [3, 4], they have limitations such as lack of explainability and the potential to learn non-causal relationships [5].
Looking ahead, multimodal data will be crucial for achieving a more comprehensive and fine-grained understanding of surgical processes, addressing the limitations of unimodal and particularly visual systems [7]. Among these modalities, acoustic sensing has emerged as a promising approach due to its radiation-free, low-cost, non-destructive, and easily integrable nature. Acoustic signals can capture time-critical events (e.g., drill breakthrough during surgical drilling [10]), provide complementary measurements to optical tracking or medical imaging, e.g., assessing press-fit during orthopedic implant insertion [9] or detect breaches during pedicle screw insertion [11], and detect characteristic acoustic events that indicate different surgical phases through ambient room microphones [8]. Despite these advances, current methods for the generation of fine-grain surgical scene representations still lack the capability to localize and correlate acoustic events within the 4D spatio-temporal context of surgery.
@@ 3.1 Experimental Setup and Data Collection
Figure 2: Overview of the experimental setup. We use our camera setup including the Ring48 microphone array, the ZED 2i RGB-D camera, and the FusionTrack 500 tracking system (shown in the left frame) to capture sequences of chiseling, drilling, and sawing (shown in the right frame).
We collected sequences of typical actions that are associated with sound emissions from Total Knee Arthoplasty using the experimental setup depicted in Figure 2. We simulate surgical sawing, drilling, and chiseling on a plastic bone model including simulated soft tissue, conducted by two trained surgical experts with real surgical tools. For drilling and sawing, we used a Arthrex DrillSaw Max 600 (Arthrex Inc., Naples, USA) surgical-grade power tool. We recorded sequences of length 20s20\>s with a frame rate of 25 fps and a sample rate of 192kHz192\>kHz from the acoustic camera for the surgical actions, resulting in $N_{chiseling}=6$, $N_{sawing}=9$, and $N_{drilling}=5$ sequences with a total length of 400s400\>s.
For the presented experiments, we use the gfai tech Ring48 (gfai tech GmbH, Berlin, Germany). The system employs a phased microphone array combined with acoustic beamforming to generate two-dimensional heatmaps of acoustic activity. An integrated optical sensor allows the acoustic activity map to be superimposed onto an optical image. To obtain a dynamic visual 3D representation of the surgical scene, we utilize a state-of-the-art RGB-D camera, namely the ZED 2i (StereoLabs, San Francisco, USA), which captures both color and depth information from the surgical scene.
@@
System Calibration and Synchronization $\mid$ All devices are calibrated and synchronized prior to the experiments. We calibrate the intrinsic parameters of the ZED 2i and Ring48 optical cameras using OpenCV and their extrinsic parameters $T_{ac}^{ftk}$ and $T_{zed}^{ftk}$ relative to the tracking system following the approach of [18], as shown in Figure 2. To ensure accurate synchronization between all devices throughout the experiments, we utilize a RocSync synchronization device [19], a hybrid clock comprising a linear and binary counter in both the RGB and IR spectrum, which enables a sub-frame accurate synchronization based on timestamps visible in all image frames. The factory-calibration of the phased microphone array is retained. At $1920\times 1080\>px$ resolution, final reprojection errors after calibration and synchronization are $2.30\pm 2.12\>px$ (Ring48) and $1.52\pm 0.84\>px$ (ZED 2i).
@@ 3.2 Multimodal Dynamic Surgical Scene Representation
We generate a multimodal, spatio-temporal scene representation based on the fusion of the acoustic information with the dynamic 3D point cloud. We compute 2D acoustic heatmaps with a fixed size of $100\times 100\>px$ via time-domain beamforming, using the professional sound analysis software NoiseImage 4.16.0 (gfai tech GmbH, Berlin, Germany). The heatmaps are computed at a predefined distance from the phased microphone array using a fixed 2D scanning grid. Band-filtering is applied between $1$ and 5kHz5\>kHz for sawing and between $1$ and 10kHz10\>kHz for drilling to obtain clear heatmaps without artifacts from ambient sound. For chiseling, no filtering is applied. The acoustic heatmaps are fused with the 3D scene representation by a simple projection operation. We project the amplitude of the acoustic heatmap, normalized to the range $[0,1]$. Future extensions may improve this step by projecting more expressive acoustic information, for example in the form of neural implicit sound descriptors.
@@ 3.3 Acoustic Event Detection
Our proposed acoustic event detection stage operates directly on the raw audio captured by the acoustic camera. We resample the audio with a sample rate of SR=16kHzSR=16\>kHz samples per second, apply a sliding window with length $L_{W}=150\>ms$ and hop length $L_{hop}=20\>ms$ to the sequence, and compute mel spectrograms with $n_{mels}=128$ mel bins from each window. For event detection, we fine-tune a pretrained AudioSpectrogramTransformer (AST) [20] in a binary classification task to predict event presence using PyTorch 2.8.0. During training we apply audio data augmentations, i.e., gain variation, gaussian noise, gain transitioning, clipping distortion, time stretching, and pitch shifting. The outputs are converted to video timestamps by detecting transitions in model predictions, as described in detail in the evaluation in Section 4. For sawing and drilling, we export multiple acoustic image frames in regular intervals after a detected event to evaluate the localization stage, as described in Section 4.
@@ 3.4 Event Localization
The location of a detected sound source is approximated within a 3D bounding box based on the computed multimodal scene representation. We localize areas with high signal amplitude within the scene using weighted clustering. Specifically, we use DBSCAN [21] with a neighborhood radius r=30.0mmr=30.0\>$m$, a minimum cluster weight of $m=200.0$, and the acoustic signal amplitudes as per-point weights. A tight bounding box is computed around the cluster with the largest total weight. Last, we limit the bounding box extents to a predefined range, based on the instrument size and the acoustic characteristics of each action. For drilling and sawing, the sound originates from the power tool surface due to motor-induced vibrations. Therefore, we limit the extents of the predicted 3D bounding box based on the known physical dimensions of the respective instruments. In contrast, for chiseling, sound emission is concentrated at the contact edge resulting from collisions between the mallet/chisel and the chisel/bone anatomy. Thus, we define a fixed-size bounding box with an edge length of 5 cm centered on the chisel’s contact edge.
@@ 4 Evaluation
Acoustic Event Detection $\mid$ Table 1 illustrates the quantitative results for the proposed event detection method described in Section 3.3. All results have been obtained using k-fold cross validation with parameters $k_{cv,chiseling}=k_{cv,drilling}=3$ and $k_{cv,sawing}=2$. We detect an event as a change in sign from $y_{pred}=0$ to $y_{pred}=1$ in the model predictions, where $y_{pred}=0$ corresponds to no event detected and $y_{pred}=1$ corresponding to a detected event. We introduce a relaxed condition, that allows for an event to be detected within a 20ms20\>ms window of $j$ frames around the ground truth event to be counted as a true positive prediction. We set this parameter as $j_{chiseling}=1$, $j_{sawing}=3$, and $j_{drilling}=10$. The hard condition only accepts predictions in the correct frame according to the ground truth. Figure 3 shows qualitative results for full example recordings.
@@
Method Precision Recall F1
@@
Chiseling $0.943\pm 0.047$ $\mathbf{0.986}\pm\mathbf{0.020}$ $0.756\pm 0.153$ $\mathbf{0.914}\pm\mathbf{0.063}$ $0.833\pm 0.101$ $\mathbf{0.948}\pm\mathbf{0.040}$
Drilling $0.000\pm 0.000$ $\mathbf{0.727}\pm\mathbf{0.196}$ $0.000\pm 0.000$ $\mathbf{0.687}\pm\mathbf{0.193}$ $0.000\pm 0.000$ $\mathbf{0.654}\pm\mathbf{0.057}$
Sawing $0.200\pm 0.000$ $\mathbf{1.000}\pm\mathbf{0.000}$ $0.200\pm 0.000$ $\mathbf{1.000}\pm\mathbf{0.000}$ $0.200\pm 0.000$ $\mathbf{1.000}\pm\mathbf{0.000}$
@@
Acoustic Event Localization $\mid$ To evaluate our event localization approach, we compute the 3D intersection over union of the predicted and ground-truth 3D bounding boxes. For drilling and sawing actions, the ground-truth bounding box is computed from the instrument mesh and tracked pose obtained from the marker-based tracking system. For chiseling sequences, we define a the ground-truth is defined as a bounding box with a side length of 5cm5\>cm centered on the chisel’s contact edge, matching the definition used during inference. The distribution of IoU scores is shown in Figure 4. In Table 2, we evaluate the localization recall rates under multiple of intersection over union thresholds ranging from $0.05$ to $0.40$. Further qualitative results are provided in the supplementary videos.
Table 2: Recall rates of the event localization. A successful event localization is defined by the 3D bounding box IoU above a defined threshold.
IoU Threshold Chiseling Drilling Sawing Average
0.05 0.79 0.95 0.91 0.88
0.10 0.76 0.86 0.73 0.77
0.20 0.68 0.39 0.31 0.46
0.40 0.37 0.00 0.01 0.13
Figure 4: Distribution of 3D bounding box IoU scores over a total of 20 recordings, grouped by surgical action.

### v2

@@ 1 Introduction
Surgical procedures are inherently dynamic, involving complex interactions among surgical staff, patients, instruments, and other medical devices. To enable the next generation of intelligent computer-aided surgery systems, capable of context-aware assistance or semi-autonomous operation, machines must possess a holistic understanding of this environment. This capability, known as surgical scene understanding, relies on creating digital representations that accurately mirror the physical reality of the operating theater.
Historically, the field has prioritized visual perception. A body of works proposed to explicitly model the surgical context based on RGB or RGB-D data, employing semantic structures such as action triplets [nwoye2020triplets] or scene graphs [oeszoy2022ssg] to capture relationships between surgeons, tools, and tissues. In contrast to explicit modeling approaches, end-to-end training has recently demonstrated strong performance in predicting high-level surgical concepts (e.g., surgical phases) from low-level observations (e.g., individual video frames) [Ozsoy_2025_CVPR, oezsoy2025egoexoregoexocentricoperatingroom]. Nevertheless, these visual-centric approaches face inherent limitations. Visual data is susceptible to occlusion, lighting variations, and often fails to capture the physical nature of tool-tissue interactions, such as the mechanical resistance of bone during sawing or the precise moment of a drill breakthrough.
Multimodal surgical perception has emerged as a promising approach to overcome the limitations of unimodal vision. Several works proposed digital twins of surgery, which integrate complementary data streams such as kinematics, inertial measurements and system events into a comprehensive digital representation of the surgical scene, effectively decoupling low-level sensing from high-level semantic reasoning [ding2024dt, Hein_2024_CVPR]. Within this multimodal landscape, audio represents a largely untapped but highly information-rich modality.
Surgical sounds provide high-frequency cues about physical interactions that are often invisible to cameras. Recently, acoustic events generated during tool–tissue interactions—such as sawing, chiseling, or drilling—have been shown to provide valuable insights into surgical tasks [seibold2022conditional, goossens2020acoustic, seibold2021realtime]. Furthermore, in our own previous work, we proposed a proof-of-concept for the localization of acoustic events in surgery based on a microphone array and acoustic beamforming in 2D image space [seibold2024ssl]. The integration of acoustic information into digital surgical scene representations introduces unique capabilities, such as high temporal resolution (sub-fps) and the ability to detect events that cannot be captured by vision alone (e.g., activity of power tools), but requires spatial and temporal integration with the visual representation to enable effective multimodal surgical scene understanding. To the best of the author’s knowledge, there is currently no comparable approach that integrates spatial sound into dynamic surgical scene representations.
@@
  • We propose a novel concept to generate 4D audio-visual representations of surgical scenes defined by a time-varying visual geometric 3D representation of a surgical scene that is enriched with localized acoustic events occurring at certain points of time or spanning over a defined period. Our approach achieves this by fusing dynamic point clouds obtained using a RGB-D camera with acoustic localization information obtained from a phased microphone array, also known as acoustic camera.
@@
The code and associated data are available under https://github.com/matthiasseibold/ssl_v2.
@@ 2 Related Work
The proposed approaches for surgical scene understanding can be categorized by the modalities used to construct the digital representation, namely purely visual approaches, acoustic sensing, and multimodal fusion. The majority of existing works relies on optical data such as RGB or RGB-D to derive both semantic and spatial information for modeling the surgical context. Early approaches focused on temporal and hierarchical modeling, using surgical process models [neumuth2017spm] to extract phases and steps from video data [padoy2019workflow]. More recently, spatial understanding has been enhanced through action triplets (instrument, action, target) [nwoye2020triplets] and surgical scene graphs, which model the operating room as a structured graph of actors and objects [oeszoy2022ssg, oeszoy2023labrador, hamoud2024stor]. While end-to-end visual models achieve strong performance on specific benchmarks [Ozsoy_2025_CVPR], they lack the ability to capture non-visual physical phenomena and are prone to learning dataset-specific visual biases [ding2024dt].
Audio has emerged as a powerful complementary modality due to its non-destructive, radiation-free, and low-cost nature, and its ability to capture high-frequency mechanical events at a sub-frame temporal resolution. Prior work has demonstrated the utility of audio for surgical phase recognition [seibold2022conditional], real-time error prevention via breakthrough detection [seibold2021realtime], or assessing the press-fit of orthopedic implants [goossens2020acoustic]. However, these methods typically treat audio as a global signal detached from the scene geometry. While our previous work introduced a proof-of-concept for 2D sound source localization in image space [seibold2024ssl], it did not account for the 3D depth and dynamic nature of the surgical environment.
The integration of multimodal data, such as vision, kinematics, language, and audio, is widely recognized as essential for next-generation computer-aided surgery [hein2017sds]. An early multimodal work by Weede et al. [weede2012workflow] fuses positional information, audio and video signals in a Bayes classifier for surgical phase recognition. In the wake of recently emerging vision-language models, a series of works fuse vision and language modalities for decision-making in surgical robotics [zargarzadeh2025decision] and workflow analysis [yuan2025learning]. Yet, current approaches generally lack the capability to localize acoustic sources. This limitation prevents the system from distinguishing between simultaneous events (e.g., two tools operating at once) or associating a specific sound with a specific instrument in a crowded scene. By projecting acoustic localization information from a phased microphone array onto dynamic point clouds, our proposed method bridges this gap, enabling the first spatially aware, multimodal representation of surgical activity.
@@ 3.1 Experimental Setup and Data Collection
Figure 2: Overview of the experimental setup. We use the Ring48 microphone array, the ZED 2i RGB-D camera, and the FusionTrack 500 tracking system (shown in the left frame) to capture sequences of chiseling, drilling, and sawing (shown in the middle frame). The right frame shows the viewpoint of the acoustic and RGB-D camera.
We collected sequences of typical actions that are associated with sound emissions from Total Knee Arthoplasty using the experimental setup depicted in Figure 2. We simulate surgical sawing, drilling, and chiseling on a plastic bone model including simulated soft tissue (SYNBONE AG, Zizers, Switzerland), conducted by two trained surgical experts with real surgical tools. For drilling and sawing, we used a Arthrex DrillSaw Max 600 (Arthrex Inc., Naples, USA) surgical-grade power tool. We recorded sequences of length 20s20\>s with a frame rate of 25 fps and a sample rate of 192kHz192\>kHz from the acoustic camera for the surgical actions, resulting in $N_{chiseling}=6$, $N_{sawing}=9$, and $N_{drilling}=5$ sequences with a total length of 400s400\>s.
For the presented experiments, we use the gfai tech Ring48 (gfai tech GmbH, Berlin, Germany). The system employs a phased microphone array with acoustic beamforming to generate two-dimensional heatmaps of acoustic activity. An integrated optical sensor allows the acoustic activity map to be superimposed onto an optical image. To obtain a dynamic visual 3D representation of the surgical scene, we utilize a state-of-the-art RGB-D camera, namely the ZED 2i (StereoLabs, San Francisco, USA), which captures both color and depth information from the surgical scene.
@@
System Calibration and Synchronization All devices are calibrated and synchronized prior to the experiments. We calibrate the intrinsic parameters of the ZED 2i and Ring48 optical cameras using OpenCV and their extrinsic parameters $T_{ac}^{ftk}$ and $T_{zed}^{ftk}$ relative to the tracking system following the approach of [hein_next-generation_2025], as shown in Figure 2. To ensure accurate synchronization between all devices throughout the experiments, we utilize a RocSync synchronization device [meyer2025rocsync], a hybrid clock comprising a linear and binary counter in both the RGB and IR spectrum, which enables a sub-frame accurate synchronization based on timestamps visible in all image frames. The factory-calibration of the phased microphone array is retained. At $1920\times 1080\>px$ resolution, final reprojection errors after calibration and synchronization are $2.30\pm 2.12\>px$ (Ring48) and $1.52\pm 0.84\>px$ (ZED 2i).
@@ 3.2 Multimodal Dynamic Surgical Scene Representation
We generate a multimodal, spatio-temporal scene representation based on the fusion of the acoustic information with the dynamic 3D point cloud. We compute 2D acoustic heatmaps with a fixed size of $100\times 100\>px$ via time-domain beamforming, using the professional sound analysis software NoiseImage 4.16.0 (gfai tech GmbH, Berlin, Germany). The heatmaps are computed at a predefined distance from the phased microphone array using a fixed 2D scanning grid. Band-filtering is applied between $1$ and 5kHz5\>kHz for sawing and between $1$ and 10kHz10\>kHz for drilling to obtain clear heatmaps without artifacts from ambient sound. For chiseling, no filtering is applied. The acoustic heatmaps are fused with the 3D scene representation by a simple projection operation. We project the amplitude of the acoustic heatmap, normalized to the range $[0,1]$.
@@ 3.3 Acoustic Event Detection
Our proposed acoustic event detection stage operates directly on the raw audio captured by the acoustic camera. We resample the audio with a sample rate of SR=16kHzSR=16\>kHz samples per second, apply a sliding window with length $L_{W}=150\>ms$ and hop length $L_{hop}=20\>ms$ to the sequence, and compute mel spectrograms with $n_{mels}=128$ mel bins from each window. For event detection, we fine-tune a pretrained AudioSpectrogramTransformer (AST) [gong21b_interspeech] in a multi-class classification task (classes = idle, chiseling, drilling, sawing) to predict event presence using PyTorch 2.8.0. During training we apply audio data augmentations, i.e., gain variation, gaussian noise, gain transitioning, clipping distortion, time stretching, and pitch shifting. The outputs are converted to video timestamps by detecting transitions in model predictions, as described in detail in the evaluation in Section 4. For sawing and drilling, we export multiple acoustic image frames in regular intervals after a detected event to evaluate the localization stage, as described in Section 4.
@@ 3.4 Event Localization
The location of a detected sound source is approximated within a 3D bounding box based on the computed multimodal scene representation. We localize areas with high signal amplitude within the scene using weighted clustering. Specifically, we use DBSCAN [ester1996density] with a neighborhood radius r=30.0mmr=30.0\>$m$, a minimum cluster weight of $\color[rgb]{0,0,0}w=400.0$}, and the acoustic signal amplitudes as per-point weights. A tight bounding box is computed around the cluster with the largest total weight. Last, we limit the side lengths of the predicted bounding box to a predefined range, based on the instrument size and the acoustic characteristics of each action. For drilling and sawing, the sound originates from the power tool surface due to motor-induced vibrations. Therefore, we limit the box side lengths to the minimum and maximum side length of the instrument’s ground-truth bounding box. In contrast, for chiseling, sound emission is concentrated at the contact edge resulting from collisions between the mallet/chisel and the chisel/bone anatomy. Thus, we define a fixed-size bounding box with an edge length of 5 cm centered on the chisel’s contact edge. Since this step is executed after the event detection, the approximate physical dimensions of the sound-emitting instruments can be assumed to be known.
@@ 4 Evaluation
Acoustic Event Detection Table 1 illustrates the quantitative results for the proposed event detection method described in Section 3.3. All results have been obtained using 3 -fold cross validation. We detect an event as a change in sign from $y_{pred}=0$ to $y_{pred}=1$ in the model predictions, where $y_{pred}=0$ corresponds to no event detected and $y_{pred}=1$ corresponding to a detected event. We introduce a relaxed condition, that allows for an event to be detected within a 20ms20\>ms window of $j$ frames around the ground truth event to be counted as a true positive prediction. We set this parameter as $j_{chiseling}=1$, $j_{sawing}=j_{drilling}=10$. The hard condition only accepts predictions in the correct frame according to the ground truth. Figure 3 shows qualitative results for full example recordings.
@@
Method Precision [%] Recall [%] F1 [%]
@@
Chiseling $0.915\pm 0.045$ $\mathbf{0.968}\pm\mathbf{0.022}$ $0.791\pm 0.134$ $\mathbf{0.953}\pm\mathbf{0.033}$ $0.845\pm 0.093$ $\mathbf{0.961}\pm\mathbf{0.026}$
Drilling $0.000\pm 0.000$ $\mathbf{0.753}\pm\mathbf{0.194}$ $0.000\pm 0.000$ $\mathbf{0.639}\pm\mathbf{0.160}$ $0.000\pm 0.000$ $\mathbf{0.651}\pm\mathbf{0.077}$
Sawing $0.611\pm 0.283$ $\mathbf{1.000}\pm\mathbf{0.000}$ $0.390\pm 0.079$ $\mathbf{0.889}\pm\mathbf{0.157}$ $0.444\pm 0.079$ $\mathbf{0.933}\pm\mathbf{0.094}$
@@
Acoustic Event Localization To evaluate our event localization approach, we compute the 3D intersection over union of the predicted and ground-truth 3D bounding boxes. For drilling and sawing actions, the ground-truth bounding box is computed from the instrument mesh and tracked pose obtained from the marker-based tracking system. For chiseling sequences, we define a the ground-truth is defined as a bounding box with a side length of 5cm5\>cm centered on the chisel’s contact edge, matching the definition used during inference. The distribution of IoU scores is shown in Figure 4 where a IoU threshold of 0.1 results in recall rates of 0.78, 0.91, 0.84, and 0.84 for chiseling, drilling, sawing, and overall average, respectively. For further insights, please refer to the supplementary material, where we evaluate the localization recall rates under multiple of intersection over union thresholds ranging from $0.05$ to $0.40$. Further qualitative results are provided in the supplementary videos.
Figure 4: Distribution of 3D bounding box IoU scores and 3D bounding box center errors over a total of 20 recordings, grouped by surgical action.
4.1 Ablation Studies
To provide further insights into the result distribution, we perform a detailed analysis of the onset error distribution for the event detection stage of the proposed method and present the results in figure 5. Furthermore, we evaluate the robustness of the event detection stage under simulated realistic OR background noise in the supplementary material. In Figure 6, we analyze the sensitivity in the choice of DBSCAN parameters, as well as the density of the input point cloud. We vary one parameter at a time in the ranges of $r\in[7.5,60]\>mm$ and $w\in[50,800]$. In addition, we analyze the robustness to the input point cloud’s density by downsampling it by a factor of $s\in[\frac{1}{64},1]$.
Last, we compare our localization stage against a naive centroid detection baseline, which computes the 3D sound source location as the weighted average of the entire point cloud, using the sound amplitude as per-point weights. To obtain a region-of-interest bounding box, an axis-aligned bounding box of the instrument’s size is centered around the estimated location. This baseline achieves a 3D bounding box IoU of $0.14\pm 0.16$ compared to $0.23\pm 0.14$ of our localization stage. The 3D bounding box center error of this baseline is $144.10\pm 121.19\>mm$ compared to $101.39\pm 89.75\>mm$ obtained by the proposed localization stage.
Figure 5: The onset error distribution for the event detection component.
Figure 6: Sensitivity analysis of the DBSCAN radius and minimum weight parameters. We additionally evaluate the effect of the point cloud density on the localization accuracy by downsampling the input point cloud, while adjusting the minimum weight accordingly. All results are reported as the average IoU on the whole dataset.